# Dependency-Aware Compilation for Surface Code Quantum Architectures

### v2

@@ 7. Discussion and Related Work
Surface code mapping and routing is an assignment of program variables to limited physical resources such that the runtime efficiency is maximized, much like the classical problem of register allocation (Cooper and Torczon, 2023; Chaitin, 1982; Poletto and Sarkar, 1999). Disjoint paths problems like the constraints on a valid gate route have also been studied in other contexts including VLSI routing (Kramer and van Leeuwen, 1984; Aggarwal et al., 2000) and all-optical networks (Aumann and Rabani, 1995; Raghavan and Upfal, 1994; Aggarwal et al., 1994). Inspired by these applications, the study of approximation algorithms for disjoint path problems is an active area with both hardness results (Andrews et al., 2005; Erlebach, 2006) and approximation algorithms on planar and grid graphs (Chuzhoy et al., 2016; Chuzhoy and Kim, 2015). However, a distinguishing feature of this particular disjoint path problem is the dependency between paths, which forces routing certain pairs in a particular order.
@@
Some recent work in compilation for surface code quantum devices has addressed similar problems to the surface code mapping and routing problem. Javadi-Abhari et al. (2017) identifies cnot contention as an important architectural design factor and develops heuristics for mapping qubits and scheduling cnot gates implemented as defect braids. Autobraid (Hua et al., 2021) builds on this work with a new algorithm for mapping that minimizes the number of large groups of cnot gates with overlapping bounding boxes and a stack-based routing algorithm. However, neither consider t gates, multiple architectures, or the constraints of lattice surgery cnot gates.
@@
For the lattice surgery setting, LSC (Watkins et al., 2023) provides a scalable framework for translation of quantum gates to lattice surgery operations, but does not optimize the mapping or routing solution. EDPC (Beverland et al., 2022a) considers a shortest-first routing algorithm, but assumes a given fixed map that places qubits in a “sparse” structure. None of the above work provides an optimal solver for the compilation problem they address. Lao et al. (2018) does solve a mapping and routing problem optimally with integer linear programming, but this formulation does not permit long-range cnot gates. An alternative proposed compilation pipeline serializes a circuit to a sequence of Pauli product rotations (Litinski, 2018; Beverland et al., 2022b), obviating the need for routing parallel operations. However, sequential Pauli compilation can lead to prohibitively high run times, converting a circuit with $k$ parallel operations to a compiled circuit of depth $k$ (Beverland et al., 2022a). LaSsynth (Tan et al., 2024) uses a sat solver to optimize lattice surgery representations of small subroutines (5-20 qubits and 10-100 operations) at a lower level of abstraction than we consider.
@@ 8. Conclusions
In this paper, we have tackled the surface code mapping and routing problem, a critical problem in compilation for emerging practical quantum computers. We developed an algorithms to deliver high-quality solutions in reasonable time via simulated annealing and dependency-awareness. In future work, we plan to extend our model and algorithms to capture a broader class of quantum architectures. For example, we could support recent designs for heterogeneous multi-chip fault-tolerant quantum architectures (Smith et al., 2022; Stein et al., 2023). The mapping and routing problem in this case is nuanced because not all cnot gates are equivalent. A cross-chip cnot gate is possible, but more costly.
@@ Appendix A Hardness of Surface Code Mapping and Routing
Given a psp instance $(J,k,t_{p})$, we can construct an equivalent scmr instance $(A,C,t_{s})$ where $A$ consists of $k$ processor units, $C$ is the union of the dependency circuit and the cycle circuit, and $t_{s}$ is computed as described above from the number of processors and the maximum number of dependencies of a job. We can translate solutions from one instance to another by executing a job in step $i$ on processor $j$ if and only if the corresponding job gadget is executed during cycle $i$ of the cycle circuit and the job circuit qubits are mapped to processor unit $j$. See LABEL:sec:hardness_details for a detailed proof.

### v3

@@ 7. Discussion and Related Work
Surface code mapping and routing is an assignment of program variables to limited physical resources such that the runtime efficiency is maximized, much like the classical problem of register allocation (Cooper and Torczon, 2023; Chaitin, 1982; Poletto and Sarkar, 1999). Disjoint paths problems like the constraints on a valid gate route have also been studied in other contexts including VLSI routing (Kramer and van Leeuwen, 1984; Aggarwal et al., 2000) and all-optical networks (Aumann and Rabani, 1995; Raghavan and Upfal, 1994; Aggarwal et al., 1994). Inspired by these applications, the study of approximation algorithms for disjoint path problems is an active area with both hardness results (Andrews et al., 2005; Erlebach, 2006) and approximation algorithms on planar and grid graphs (Chuzhoy et al., 2016; Chuzhoy and Kim, 2015). However, a distinguishing feature of this particular disjoint path problem is the dependency between paths, forcing the routing of certain pairs in a particular order.
@@
Some recent work in compilation for surface code quantum devices has addressed similar problems to the surface code mapping and routing problem. Javadi-Abhari et al. (2017) identifies cnot contention as an important architectural design factor and develops heuristics for mapping qubits and scheduling cnot gates implemented as defect braids. Autobraid (Hua et al., 2021) builds on this work with a new algorithm for mapping that minimizes the number of large groups of cnot gates with overlapping bounding boxes and a stack-based routing algorithm. However, neither considers t gates, multiple architectures, or the constraints of lattice surgery cnot gates.
@@
For the lattice surgery setting, LSC (Watkins et al., 2024) provides a scalable framework for translation of quantum gates to lattice surgery operations, but does not optimize the mapping or routing solution. EDPC (Beverland et al., 2022a) considers a shortest-first routing algorithm, but assumes a given fixed map that places qubits in a “sparse” structure. None of the above work provides an optimal solver for the compilation problem they address. Lao et al. (2018) does solve a mapping and routing problem optimally with integer linear programming, but this formulation does not permit long-range cnot gates. LaSsynth (Tan et al., 2024) uses a sat solver to optimize lattice surgery representations of small subroutines (5-20 qubits and 10-100 operations) at a lower level of abstraction.
An alternative proposed compilation pipeline serializes a circuit to a sequence of Pauli product rotations (Litinski, 2019; Beverland et al., 2022b). In this setting, TopQAD (Silva et al., 2024) applies an algorithm similar to LSC: performing no mapping optimization and routing gates in a random order. The only difference between this greedy routing algorithm and dascot-randroute is in the specifics of the routed operations. Pauli product rotations are implemented via routing trees with an interacting qubit at each leaf, rather than routing paths with an interacting qubit at each endpoint. However, sequential Pauli compilation can lead to prohibitively high run times, converting a circuit with $k$ parallel operations to a compiled circuit of depth $k$ (Beverland et al., 2022a).
@@ 8. Conclusions
In this paper, we have tackled the surface code mapping and routing problem, a critical problem in compilation for emerging practical quantum computers. We developed an algorithm to deliver high-quality solutions in reasonable time via simulated annealing and dependency-awareness. In future work, we plan to extend our model and algorithms to capture a broader class of quantum architectures. For example, we could support recent designs for heterogeneous multi-chip fault-tolerant quantum architectures (Smith et al., 2022; Stein et al., 2023). The mapping and routing problem in this case is nuanced because not all cnot gates are equivalent. A cross-chip cnot gate is possible, but more costly.
Data-Availability Statement
The software and benchmark suite that supports our evaluation in Section 6 is available in a public archive (Molavi et al., 2025). This software artifact includes a Docker image packaging the source code for dascot and the baseline algorithms, along with scripts for reproducing our empirical results and generating plots matching those included in the paper.
Acknowledgments
We thank the anonymous reviewers for their insightful feedback and suggestions. We are grateful to the CHTC team for maintaining the computing resources which enabled our empirical evaluation (Center for High Throughput Computing, 2006). This work is supported by nsf grants #1652140 and #2212232 and awards from Meta and Amazon. This research is also partially supported by the OVCRGE at the University of Wisconsin-Madison with funding from the Wisconsin Alumni Research Foundation.
@@ Appendix A Hardness of Surface Code Mapping and Routing
Given a psp instance $(J,k,t_{p})$, we can construct an equivalent scmr instance $(A,C,t_{s})$ where $A$ consists of $k$ processor units, $C$ is the union of the dependency circuit and the cycle circuit, and $t_{s}$ is computed as described above from the number of processors and the maximum number of dependencies of a job. We can translate solutions from one instance to another by executing a job in step $i$ on processor $j$ if and only if the corresponding job gadget is executed during cycle $i$ of the cycle circuit and the job circuit qubits are mapped to processor unit $j$. The rest of this section elaborates on each of these proof steps.